OpenStudy (anonymous):

how do i integrate : integral from 0 to 1 of pi[4tan(pi/3*y)]^2

OpenStudy (anonymous):

16pi tan^2(y pi/3)= 16 pi (sec^2(y pi / 3)-1) then you can go from there, i think amistre and i did this one for you earlier

OpenStudy (anonymous):

u = pi / 3 y 3 / pi du= dy INT 3 / pi (16pi)sec^2(u)-1 dy 48(tan(u)-y)+C since u=pi/3y 48((tan(pi / 3 y)-y)+C

OpenStudy (anonymous):

then from [0,1] since tan(0)=0 and y=0 just use the top bound to evaluate the integral 48(sqrt(3)-1)